9. Look up the current price and calculate the P/E ratio for the following companies:

Company Ticker EPS Current Price P/E Ratio

Johnson & Johnson JNJ $0.45 $

Walmart, Inc. WMT $3.00 $

Pfizer, Inc. PFE $3.60 $

Alphabet, Inc. GOOGL $23.62 $

10. Dividend discount model: Using the following data provided, calculate the current price of these stocks assuming an investor required rate of return of 7%. Company Next year’s dividend Next year’s expected price Current value Alfred Publishing $0.45 $22.00 Aldo Food Market $1.75 $48.00 ABC Warehouse $2.80 $72.00 Anderson Windows $20.00 $1000.00

11. Constant-growth dividend discount model: Using the following data provided, calculate the current price of these stocks assuming an investor required rate of return of 9%. Company Current year’s dividend Dividend growth rate Current price Cabo Technology $0.90 5.0% Cellular Systems $3.25 3.0% Candida Consultants $5.38 6.0% Coliseum Theaters $12.75 7.0%
Answer #1

9)

Company Johnson & Johnson Walmart Pfizer Alphabet Ticker EPS Current price P/E ratio 130.69 290.42 32.58 11.81 46.88 JNJ 0.45 3.00 3.60 GOOGL 23.62 WMT PFE 97.73 42.53 1,107.30 Compan 2 Johnson & Johnson 3 Walmart 4 Pfizer 5 Alphabet 6. Ticker JNJ WMT PFE GOOGL EPS 0.45 Current price 130.69 97.73 42.53 1107.3 P/E ratio D2/C2 D3/C3 D4/C4 D5/C5 3.6 23.62

10)curent price=(dividend+next year price)/(1+required return)
Alfred Publishing=(0.45+22)/(1+7%)=20.98
Aldo Food Market =(1.75+48)/(1+7%)=46.50
ABC Warehouse =(2.8+72)/(1+7%)=69.91
Anderson Windows =(20+1000)/(1+7%)=953.27

11)current price=current dividend*(1+growth)/(required return-growth)
Cabo Technology=(0.9*(1+5%))/(9%-5%)=23.63
Cellular Systems =(3.25*(1+3%))/(9%-3%)=55.79
Candida Consultants=(5.38*(1+6%))/(9%-6%)=190.09
Coliseum Theaters=(12.75*(1+7%))/(9%-7%)=682.13